OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 1.18x + 4.34 = 8.67 A. 11.0 B. 3.8 C. 3.7 D. 3.6

OpenStudy (anonymous):

i think ts 3.6 but im not really sure

OpenStudy (anonymous):

1.18x+4.34=8.67 -4.34 -4.34 1.18x=4.33 1.18 1.18 x=3.66 Answer is 3.7

OpenStudy (anonymous):

3.1x + 5.2 = 9.9

OpenStudy (anonymous):

Estimate the solution to the equation by first rounding the constants and coefficients to the nearest integer.

OpenStudy (anonymous):

Follow how I did it above:) If you need help through the steps I can help

OpenStudy (anonymous):

A. 1.3 B. 1.5 C. 1.6 D. 1.7

OpenStudy (anonymous):

i said it was 1.7

OpenStudy (anonymous):

Correct